OpenStudy (anonymous):

evaluate log base 6 of the fifth root of 36

OpenStudy (whpalmer4):

Hint: \[\sqrt[5]{36} = 36^{\frac{1}{5}}\]

OpenStudy (anonymous):

okay, but I'm kind of stuck because 36 is the square of 6 and I don't know how I should evaluate it with the 1/5 if that makes any sense at all

OpenStudy (whpalmer4):

okay, if 6 is the square root of 36, then \(36 = 6^2\) and we can rewrite our problem as \[\sqrt[5]{36} = (6^2)^{\frac{1}{5}}\]Now, do you remember how to simplify \[(a^b)^c\]?

OpenStudy (anonymous):

um that would be a^b x c, right?

OpenStudy (whpalmer4):

\[(a^b)^c = a^{b*c}\] so that gives us \[6^{\frac{2}{5}}\]How would we find \[\log_{6}6^{\frac{2}{5}}\]

OpenStudy (anonymous):

oh my gosh thank you so much that makes a lot more sense

OpenStudy (whpalmer4):

hint: \[\log_{b}b^n=n\]

OpenStudy (anonymous):

okay, so the solution would be 2/5 thank you very much

OpenStudy (whpalmer4):

exactly!

OpenStudy (whpalmer4):

one little quibble from earlier: you wrote "that would be a^b x c, right?" I know what you meant, but really you should have written that as a^(b x c) because the operator precedence is that exponentiation comes before multiplication, so what you actually wrote is c x a^b

OpenStudy (whpalmer4):

just trying to avoid a potential wrong answer for you down the road :-)

OpenStudy (anonymous):

okay, okay :) thanks for getting what i mean, despite my bad form